# Federated Learning and Differential Privacy Techniques on Multi-hospital Population-scale Electrocardiogram Data

### v1

@@ 3.3. Learning Algorithm
For the Deep Learning (DL) model, we implemented a CNN based on the residual neural network architecture, consisting of a convolutional layer, 4 residual blocks with 2 convolutional layers per block, followed by a dense layer (total of 8,922,644 model parameters) (Sun2023-lp, ). We used batch normalization (Ioffe2015-yu, ), ReLU (Agarap2018-uw, ), and dropout after each convolutional layer. Each ECG instance was loaded as a 12×5120 numeric matrix. Additional features such as age and sex were passed to a 10-hidden-unit layer and concatenated with the dense layer, and finally passed to a softmax layer to produce the outputs. Binary cross-entropy was used as the loss function with the initial learning rate of $10^{-3}$, Adam optimizer (Kingma2014-qn, ), ReLU activation function, kernel size of 16, batch size of 512, and dropout rate of 0.2 with other hyper-parameters set to default. Models were learnt for a maximum of 50 epochs. The learning rate was reduced to $10^{-5}$ if there was no improvement in tuning loss for 7 consecutive epochs, and the learning process was stopped if loss in the tuning set did not reduce for 9 epochs. We trained all our models on the NVIDIA Driver version 418.88 with 8 Tesla V100-SXM2 GPUs and 32 GB of RAM per GPU. Each DL model took approximately 30 min per epoch to train.
@@ 3.4.1. Federated Averaging (FedAvg)
We implemented the Federated Averaging (FedAvg) algorithm (pmlr-v54-mcmahan17a, ), a decentralized machine learning paradigm designed for training models across various hospitals while ensuring data privacy. In the FedAvg framework, each participating hospital autonomously conducts model training on its locally stored ECG data. Instead of transmitting raw data to a central server, only the model updates, typically in the form of gradients, are communicated. The central server aggregates these updates and computes a global model through parameter averaging. Subsequently, the resulting global model is disseminated to the participating hospitals, initiating a recurrent iterative process. FedAvg is particularly beneficial in scenarios marked by distributed ECG data across multiple sites, enabling collaborative model training without explicit data exchange.
@@ 3.4.2. Differentially Private Stochastic Gradient Descent (DP-SGD)
To implement differential privacy during model training, we have employed the Differential Privacy Stochastic Gradient Descent (DP-SGD) algorithm (Abadi2016-fu, ). This advanced optimization technique is specifically tailored for machine learning models, primarily focusing on integrating differential privacy guarantees. DP-SGD extends the traditional Stochastic Gradient Descent (SGD) optimization algorithm by introducing mechanisms that introduce controlled noise into gradient computations. This approach is particularly effective in mitigating the risk of extracting sensitive information, such as patient details, from the ECG data. The intentional addition of noise ensures that the model’s updates maintain differential privacy, whereby the inclusion or exclusion of any single ECG has a negligible impact on the overall model parameters. DP-SGD strikes a balance between optimizing model performance and upholding the privacy of patient information within the ECG data. We have incorporated DP-SGD into our framework utilizing the Opacus library (opacus, ), which employs the Rényi Differential Privacy (8049725, ; Mironov2019-we, ) accountant as its mechanism for tracking privacy.
@@ 3.5. Evaluation
We split the ECG dataset into different groups for training the models and evaluating models. We randomly split the ECGs by disjoint patients into the development set (60%) and holdout set (40%). In the development set, we split into 80% for training and 20% for tuning the model’s performance during development. The proportions were maintained approximately within each hospital site for the previously mentioned split. The holdout set was kept aside to independently test the model. To mitigate the bias, we made sure that ECGs from the same patient were not used in both the development and holdout sets. The metric we reported was based on this holdout set, particularly the area under the receiver operating characteristic curve (AUROC) (Bradley1997-fv, ), also known as the C-index. For each of the 10 ICD codes corresponding to specific medical diagnoses, we conducted individual calculations of the AUROC values. Subsequently, we computed the macro AUROC score encompassing all diseases. This iterative process was repeated 1000 times through repeated random sampling, involving 10% of the testing data on each occasion. Following collecting macro AUROCs from these iterations, we calculated the mean of the 1000 macro AUROCs and determined the 95% confidence intervals using the percentile method. We use these 95% confidence intervals to discern statistical significance in our results.
@@ 4. Results
In this section, we present the results from our analysis of patient characteristics and the experiments conducted for our diagnosis task using FL and DP. Table 2 and Table 3 present the distribution of age, sex, common comorbidities and diagnostic classes in the study cohort across the hospitals, highlighting their differences. Hospitals 1 and 2 consistently exhibit the highest prevalence percentages among all participating healthcare facilities for most classification labels. Moreover, the distribution of gender and age across all hospitals is notably varied. Hospitals 4, 5, and 7 exhibit a higher median age range of 71-72 years, which is 5 to 8 years greater compared to Hospitals 1, 2, and 4, where the median age ranges from 64 to 66 years. Similarly, Hospitals 1 and 2 have a relatively higher percentage of men (58-59%) compared to Hospital 4, which maintains an equal gender distribution with 50% men and 50% women. All comorbidities and diagnostic classes also showed significant variations in their prevalence rates across the hospitals. Noteworthy variations are observed in the prevalence rates of specific comorbidities, with Peripheral Vascular Disease exhibiting a significantly higher prevalence in Hospital 3 (9.50%) compared to other hospitals (with a maximum of 1.48%). Conversely, the prevalence of Cancer, Dementia, Heart Failure (I50.0), Paroxysmal Atrial Fibrillation (I48.0), and Myocardial Infarction is conspicuously lower in Hospital 6 in comparison to its counterparts.
4.1. Results from Federated Learning
In this section, we examine the outcomes resulting from the FL model training process on our dataset. Note that, from this point onward, we refer to hospitals as sites in this section. Initially, we employed a conventional approach, where the entire development dataset (training and tuning ECGs from all the hospitals) was utilized to train a single ECG classification model. This model will serve as our benchmark for evaluation, as we aim to attain comparable performance using an FL setup. Subsequently, we implemented a DL model with a similar architecture within the framework of FL. We then conducted a comparative analysis of the results obtained from both models using the holdout dataset (holdout ECGs from each of the hospitals, as well as the entire holdout set). For reference, we will denote the first approach as the ”standard approach” method and the second as the ”FL approach.”
@@ 4.1. Results from Federated Learning
As depicted in Table 4 below, we conducted a comparative analysis of models trained using 2 distinct methodologies. The first row outlines the model’s performance when trained under the conventional approach, which involves centralizing all data for model training. Subsequently, the second row illustrates the model’s performance when trained using the FL approach. We observe that specific sites, such as Site 7, exhibit low performance even when employing the Standard approach. It is also noteworthy that a marginal reduction in performance is observed with the adoption of the FL approach as opposed to the conventional data aggregation technique. Furthermore, the performance difference between the standard approach model and the FL approach model is statistically significant only for Site 1’s test data, while the test performance on data from all other hospitals is not statistically different. Nonetheless, it is crucial to acknowledge that the FL method brings substantial benefits in strengthening data security and ensuring privacy as described by Yin et al. (Yin2022-st, ). The results are presented graphically in Figure 4, given by the data provided in Table 4. It is essential to highlight that Figure 4 represents the mean of the Macro AUROCs derived from the testing dataset, accompanied by their corresponding 95% confidence intervals. The rationale behind choosing macro AUROC over label-wise performance lies in its alignment with the trends observed in the class-wise AUROC scores (refer to Figures 7 and 8 in the Appendix). Utilizing a single value, the Macro AUROC simplifies summarizing the overall performance, avoiding the complexity of interpreting 10 distinct values. Moreover, we have observed from Table 3 and Figures 7 and 8 that there is no discernible relation between the prevalence of diseases and their corresponding AUROC scores.
@@
In the subsequent experiment, we conduct both intra-site and inter-site testing for our ECG classification model without utilizing FL. This experiment is designed to encourage the adoption of FL for training classification models. Figure 5 depicts the schematic representation of this experiment. A classification model is trained without FL on Hospital A and subsequently tested on data from Hospital B’s testing dataset for inter-site testing. Conversely, for intra-site testing, the model is trained and tested on Hospital A’s training and testing datasets, respectively. Table 5 offers a comprehensive overview of the performance of models trained utilizing data from individual sites. For instance, let us consider the cell denoted as (site-1, site-1) in the table. This specific cell represents the model’s performance when trained using data exclusively from site-1 and subsequently tested on the test data originating from the same site. The last column of the table contains the model’s overall performance on the entire test dataset. Similarly, the final row in the table presents the model’s performance across various sites when trained on the entire training dataset. This row is the same as the standard approach in the previous experiment (Table 4).
A conspicuous pattern emerges from the presented table. The models trained on sites 6 and 7 demonstrate inferior performance compared to all models trained on data from the first 5 sites when tested on any of the 7 test sets. This may be attributed to the first 5 sites possessing significantly more number of ECG samples compared to sites 6 and 7. This does not imply that site-1 consistently achieves the best performance across all test sets, as the model trained on site-2 data exhibits superior performance in 4 out of the 7 test sets. Furthermore, the observed trend reveals a direct and proportional relationship between the model’s performance and the origin of its training data with the exception of Sites 6 and 7. A model trained on site-A exhibits performance that matches those trained on other sites when tested on site-A data. For instance, a model trained on site-5 training data performs noticably worse than other models when tested on datasets other than the site-5 testing dataset. However, when tested on site-5 data, its performance is marginally better to any of the other models. However, a diminishing performance trend is evident in other sites as their respective training dataset sizes decrease. When comparing the performance of models trained individually for each site with those trained using the standard approach with the complete train dataset, a clear pattern emerges. In the majority of instances, the standard approach model surpasses its individually trained counterparts. This trend underscores the significant advantages of adopting the FL approach, as it consistently enhances the model’s performance across most cases. Particularly, hospitals with smaller sample sizes experience a marked improvement in their model performance, in stark contrast to hospitals with larger sample sizes, which exhibit a more modest impact.
4.2. Results from Differential Privacy
@@
In our experiment with DP, we investigate the impact of incorporating DP on our model’s performance. Our attention shifts to the model training process, employing the FL approach and integrating DP at the site level. During the local training phase within this FL context, we implemented DP-SGD for parameter updates. As noted in Section 1.2, the parameter $\epsilon$ in DP measures the privacy guarantees provided by the algorithm. A smaller value of $\epsilon$ signifies a stronger privacy guarantee, implying that less information about a specific individual can be inferred from the output. It fine-tunes the balance between data utility and individual privacy when implementing DP. Thus, as privacy in DP is quantified using $\epsilon$, we varied this parameter to observe its influence on model performance.
During our testing with different $\epsilon$ values in the DP-SGD technique, we observed a noticeable pattern (see Figure 6). It became evident that the model’s performance exhibited a diminishing trend as we reduced the $\epsilon$ value. The model performances approached AUROC levels akin to random guessing, with values close to 0.5, for $\epsilon$ values below 1. $\epsilon$ serves as a pivotal element in DP, as it governs the extent of noise added to the gradient updates during training. As we lower $\epsilon$, a more stringent level of privacy protection is achieved, accompanied by increased noise injection. However, we have observed from our experiments that this heightened privacy protection comes at the cost of reduced model performance.
@@ 5. Discussion
Our analytical framework comprises 2 distinct components: FL and DP. In the FL framework, we benchmarked our model performance against standard models devoid of the FL technique. Our findings indicate that FL model performance registers a negligible decrease (¡ 2%) across all sites, concurrently ensuring robust data security and privacy. During inter-site testing, we also observed advantageous outcomes for hospitals with limited data under the FL paradigm. This advantage may be due to the potential inclusion of rare disease occurrences in datasets from hospitals with more extensive ECG collections. We also note that models trained using the standard approach on sites with many ECGs (e.g., Site 1 and Site 2) demonstrate good performance when tested across all sites, including those with fewer ECGs. Furthermore, we observe a direct correlation between the origin of training data and model performance. Models perform well when trained and tested on the same site, as displayed by a model trained on Site 5 data and tested on Site 5 data, exhibiting comparable performance to models tested on other sites for the corresponding testing data. Moreover, our model shows increased robustness within the FL framework, owing to its training across disparate datasets with varying demographics and prevalence rates. This adaptability enhances the model’s generalization capacity, fostering resilience to variations in data.
In the DP experiment, the DP-SGD algorithm was employed, revealing a discernible correlation between the value of $\epsilon$ and the model performance. As $\epsilon$ increases, indicating a decrease in privacy, a concurrent increase in model performance is observed—an anticipated outcome. The trade-off between privacy and utility becomes more pronounced with diminishing $\epsilon$ values. While the augmented noise levels enhance privacy safeguards, they also diminish the model’s ability to discern subtle patterns and features in the data. Consequently, the model’s performance deteriorates when $\epsilon$ values are reduced, manifesting as reduced accuracy and predictive capability. This observation underscores the fundamental trade-off that must be considered in applications of differential privacy: increasing privacy constraints may augment utility at the potential expense of privacy, whereas augmenting privacy protection can lead to compromised model performance. Therefore, the selection of epsilon values should align with the project’s specific privacy requirements and performance objectives. Ponomareva et al. (Ponomareva2023-me, ) posits that maintaining an epsilon value below 10 provides a substantiated privacy assurance. The assertion is supported by empirical evidence derived from real-world datasets and machine-learning models.
@@
Recent research has suggested alternative methodologies for conducting federated averaging on non-independent and identically distributed (non-IID) data (Tenison2022-uh, ; Hong2022-yd, ; Xiao2020-sa, ). Despite exploring these advanced models beyond FedAvg in domains such as healthcare, we used FedAvg as a baseline model for this study. This decision is based on its role in evaluating the initial feasibility of retaining the FL approach, recognizing the preliminary nature of our investigation into this specific topic. Our choice of labels in the methods section was driven by thoughtful consideration of ICD-10 codes associated with cardiovascular and metabolic conditions of clinical importance. The decision-making process was informed by ECG-based predictability of selected cardiovascular ICD-10 codes (Sun2022-nx, ), and sufficient representation across various hospital sites. However, we acknowledge the representative nature of our label selection and highlight the need for future experiments to explore if similar trends extend to other diagnostic categories. The use of ECG machines from a single manufacturer to capture all ECG data in this study presents a potential limitation. Employing ECG equipment from a singular manufacturer restricts the diversity of instrumentation and may limit the generalizability and external validity of the study’s findings. However, the examination of ECG cohorts reveal notable variations in age, sex and disease prevalence rates across the 7 hospitals despite their geographical proximity and single healthcare provider (Table 2 and Table 3). These observed variations in clinical and demographic distributions among hospitals underscore the importance for adopting FL approaches for developing robust and equitable prediction models.
@@
The imperative for rigorous adherence to regulatory frameworks in healthcare, exemplified by the Health Insurance Portability and Accountability Act (HIPAA) in the United States and the General Data Protection Regulation (GDPR) in the European Union, underscores the importance of implementing FL and DP models for diagnosis tasks. Given the sensitive nature of health data and the potential for significant privacy breaches, compliance with such regulations is not merely a legal obligation but a fundamental ethical consideration. Against this backdrop, FL emerges as a pivotal research area with its inherent focus on decentralized model training and DP. This study seeks to explore and contribute to the intersection of FL and DP, aiming to enhance ECG ML models’ robustness and privacy assurances in compliance with prevailing legal frameworks. DP-SGD offers several noteworthy advantages in ML and data privacy. First and foremost, its incorporation of gradient clipping and noise addition as defense mechanisms against data poisoning significantly enhances the resilience of ML models. By constraining gradients and introducing controlled noise, DP-SGD mitigates the risk of adversarial influence on the training process, ultimately leading to more reliable and robust models. Moreover, DP-SGD’s ability to safeguard against privacy breaches, including membership inference attacks, through strategic noise injection is a crucial feature. The strategic noise injection ensures the privacy of individuals’ data and hinders malicious actors from inferring sensitive information, thereby reinforcing privacy protections. These advantages position DP-SGD as a valuable tool in addressing the twin challenges of data poisoning and privacy breaches in machine learning. However, it is essential to acknowledge certain limitations associated with DP-SGD. One prominent concern is the trade-off between privacy and utility. The introduction of noise into the gradient updates, although essential for privacy protection, can reduce the model’s overall accuracy and performance. Striking the right balance between preserving privacy and maintaining a high level of utility remains a complex and ongoing challenge. One potential approach for determining the optimal value of $\epsilon$ in the context of DP might be employing Membership Inference Attacks (7958568, ) to evaluate the effectiveness of the privacy measures. Additionally, the computational overhead incurred by DP-SGD is non-negligible. The process of computing and adding noise to gradients can significantly increase the training time and resource requirements, which can be a practical concern in applications where real-time or high-throughput processing is critical. Therefore, practitioners and researchers must carefully consider these trade-offs and make informed decisions when implementing DP-SGD in their machine learning workflows. Nonetheless, the benefits of DP-SGD in terms of privacy preservation and robustness make it a compelling approach, with its limitations being areas for continued refinement and innovation in privacy-preserving machine learning.

### v2

@@ 3.3. Learning Algorithm
For the DL model, we implemented a CNN based on the residual neural network architecture, consisting of a convolutional layer, 4 residual blocks with 2 convolutional layers per block, followed by a dense layer (total of 8,922,644 model parameters) (Sun2023-lp, ). We used batch normalization (Ioffe2015-yu, ), ReLU (Agarap2018-uw, ), and dropout after each convolutional layer. Each ECG instance was loaded as a 12×5120 numeric matrix. Additional features such as age and sex were passed to a 10-hidden-unit layer and concatenated with the dense layer, and finally passed to a softmax layer to produce the outputs. Binary cross-entropy was used as the loss function with the initial learning rate of $10^{-3}$, Adam optimizer (Kingma2014-qn, ), ReLU activation function, kernel size of 16, batch size of 512, and dropout rate of 0.2 with other hyper-parameters set to default. Models were learnt for a maximum of 50 epochs. The learning rate was reduced to $10^{-5}$ if there was no improvement in tuning loss for 7 consecutive epochs, and the learning process was stopped if loss in the tuning set did not reduce for 9 epochs. We trained all our models on the NVIDIA Driver version 418.88 with 8 Tesla V100-SXM2 GPUs and 32 GB of RAM per GPU. Each DL model took approximately 30 minutes per epoch to train.
@@ 3.4.1. Federated Averaging (FedAvg)
We implemented the Federated Averaging (FedAvg) algorithm (pmlr-v54-mcmahan17a, ), a decentralized machine learning paradigm designed for training models across various hospitals while ensuring data privacy. In the FedAvg framework, each participating hospital autonomously conducts model training on its locally stored ECG data. Instead of transmitting raw data to a central server, only the model updates, in the form of gradients, are communicated. The central server aggregates these updates and computes a global model through parameter averaging. Subsequently, the resulting global model is disseminated to the participating hospitals, initiating a recurrent iterative process. FedAvg is particularly beneficial in scenarios marked by distributed ECG data across multiple sites, enabling collaborative model training without explicit data exchange.
@@ 3.4.2. Differentially Private Stochastic Gradient Descent (DP-SGD)
To implement differential privacy during model training, we employed the Differential Privacy Stochastic Gradient Descent (DP-SGD) algorithm (Abadi2016-fu, ). This optimization technique focuses on integrating differential privacy guarantees. DP-SGD extends the traditional Stochastic Gradient Descent (SGD) optimization algorithm by introducing controlled noise into gradient computations. This approach is particularly effective in mitigating the risk of extracting sensitive personal information from the ECG data. The intentional addition of noise ensures that the model’s updates maintain differential privacy, whereby the inclusion or exclusion of any single ECG has a negligible impact on the overall model parameters. DP-SGD strikes a balance between optimizing model performance and upholding the privacy of patient information within the ECG data. We have incorporated DP-SGD into our framework utilizing the Opacus library (opacus, ), which employs the Rényi Differential Privacy (8049725, ) accountant as its mechanism for tracking privacy.
@@ 3.5. Evaluation
We randomly assigned the ECGs into the development set (60%) and holdout set (40%). In the development set, we used 80% for training and 20% for tuning the model’s performance. The proportions were maintained approximately within each hospital site (Table 1). The holdout set was kept aside to independently test the model. To mitigate the bias, we made sure that ECGs from the same patient were not used in both the development and holdout sets. For each of the 10 ICD code labels corresponding to specific medical diagnoses, we calculated the area under the receiver operating characteristic curve (AUROC) (Bradley1997-fv, ) values based on the holdout set. Subsequently, we computed the macro AUROC score encompassing all diseases. This process was repeated 1000 times through repeated random sampling, involving 10% of the holdout data on each occasion. Then, we calculated the mean of the 1000 macro AUROCs and determined the 95% confidence intervals using the percentile method. We used these 95% confidence intervals to discern statistical significance in our results.
@@ 4. Results
4.1. Characteristics of Multi-hospital Cohorts
Table 2 presents the distribution of age, sex, common comorbidities and diagnostic classes in the study cohort across the hospitals, highlighting their differences. Hospitals 1 and 2 consistently exhibited the highest prevalences among all participating healthcare facilities for most classification labels. The distribution of gender and age across all hospitals is notably varied. Hospitals 4, 5, and 7 showed a higher median age range of 71-72 years, which is 5 to 8 years greater compared to Hospitals 1, 2, and 4, where the median age ranged from 64 to 66 years. Similarly, Hospitals 1 and 2 had a relatively higher percentage of men (58-59%) compared to Hospital 4, which maintained approximately equal gender distribution. All comorbidities and diagnostic classes also showed significant variations in their prevalence rates across the hospitals. Noteworthy variations are observed in the prevalence rates of specific comorbidities, with Peripheral Vascular Disease exhibiting a significantly higher prevalence in Hospital 3 (9.50%) compared to other hospitals (with a maximum of 1.48%). Conversely, the prevalence of Cancer, Dementia, Heart Failure (I50.0), Paroxysmal Atrial Fibrillation (I48.0), and Myocardial Infarction was conspicuously lower in Hospital 6 compared to others.
4.2. Federated Learning
Initially, we employed a conventional approach, where the entire development dataset (combining training and tuning ECGs from all the hospitals) was utilized to train a single ECG classification model. This model served as our benchmark for evaluation, as we aim to attain comparable performance using an FL setup. Subsequently, we implemented a DL model with a similar architecture within the framework of FL. We then conducted a comparative analysis of the results obtained from both models (holdout ECGs from each of the hospitals, as well as the entire holdout set). For reference, we will denote the first approach as the ”pooled approach” method and the second as the ”FL approach” (Table 3, Figure 4). We observe that specific sites, such as Site 7, exhibit low performance even with the Pooled approach. A marginal reduction in performance was observed with FL approach as opposed to the conventional data aggregation technique. However, the performance difference between the Pooled approach model and the FL approach model was statistically significant only for Site 1’s test data, while the test performance on data from all other hospitals was not statistically different. When we examined the class-wise AUROCs, we did not observe any discernible correlation between the prevalence of diseases (i.e., positive rate of labels) and their corresponding AUROC scores (Table 2, Appendix: Figures 7 and 8).
@@ 4.2. Federated Learning
In the subsequent experiment, we conducted both intra-site and inter-site testing for our ECG classification model without utilizing FL (Figure 5). This experiment was designed to explore how individual sites varied in their performance benefits from the FL framework, compared to local model development. Table 4 shows performance of models trained utilizing data from individual sites.
The models trained on sites 6 and 7 demonstrated inferior performance compared to all models trained on data from the first 5 sites and tested on any of the 7 test sets. This may be attributed to the first 5 sites having significantly more number of ECG samples compared to sites 6 and 7. However, site-1 did not achieve the best performance across all test sets consistently, as the model trained on site-2 data exhibits superior performance in 4 out of the 7 test sets. Moreover, a diminishing performance trend was observed in other sites as their respective training dataset sizes decreased. Furthermore, we found that a model trained on a particular site exhibits performance that matches those trained on other sites when tested on the same site data, except for Sites 6 and 7. In other words, intra-site testing performs better than inter-site testing. For instance, a model trained on site-5 training data performs noticeably worse than other models when tested on datasets other than the site-5 testing dataset. However, when tested on site-5 data, its performance is marginally better than any of the other models. This could be attributed to covariate and distributional shifts across the sites. When comparing the performance of models trained individually for each site with those trained using the pooled approach with the complete train dataset, we found that both the pooled and FL approach models clearly surpassed their individually trained counterparts. This trend underscores the significant advantages of adopting the FL approach, as it consistently enhances the model’s performance across most sites. Particularly, hospitals with smaller sample sizes showed a marked improvement in their model performance, while hospitals with larger sample sizes exhibited a more modest impact.
4.3. Differential Privacy
@@
During the local training phase within FL context, we implemented DP-SGD for parameter updates at the site level. As noted in Section 1.2, the parameter $\epsilon$ in DP indicates the privacy guarantees provided by the algorithm. A smaller value of $\epsilon$ signifies a stronger privacy guarantee, implying that less information about a specific individual can be inferred from the output. As privacy is quantified using $\epsilon$, we varied this parameter to observe its influence on model performance. The model’s performance exhibited a similar diminishing trend for all of the seven sites as we reduced the $\epsilon$ value (Figure 6). The model performances approached AUROC levels akin to random guessing, with values close to 0.5, for $\epsilon$ values below 1. Overall, we observed that heightened privacy protection comes at the cost of reduced model performance consistently across all sites.
@@ 5. Discussion
The imperative for rigorous adherence to regulatory frameworks in healthcare, exemplified by the Health Insurance Portability and Accountability Act (HIPAA) in the United States and the General Data Protection Regulation (GDPR) in the European Union, underscores the importance of implementing FL and DP models for diagnosis tasks. Given the sensitive nature of health data and the potential for significant privacy breaches, compliance with such regulations is not merely a legal obligation but a fundamental ethical consideration. Against this backdrop, FL emerges as a pivotal research area with its inherent focus on decentralized model training and DP. This study contributes some novel findings to this end, aiming to enhance ECG ML models’ robustness and privacy assurances in compliance with prevailing legal frameworks.
@@
Our analytical framework comprises 2 distinct components: FL and DP. In the FL framework, we benchmarked our model performance against pooled models devoid of the FL technique. Our findings indicate that FL model performance registers a negligible decrease (¡ 2%) across all sites compared to the pooled model trained using data from all the hospitals. During inter-site testing, we observed advantageous outcomes for hospitals with limited data under the FL paradigm. This advantage may be due to the potential inclusion of rare disease occurrences in datasets from hospitals with more extensive ECG datasets. We also note that models trained using the pooled approach on sites with many ECGs (e.g., Site 1 and Site 2) demonstrate good performance when tested across all sites, including those with fewer ECGs. Furthermore, we observe a direct correlation between the origin of training data and model performance. Models perform well when trained and tested on the same site, as displayed by a model trained on Site 5 and tested on Site 5, exhibiting comparable performance to models tested on other sites for the corresponding testing data. Moreover, our model shows increased robustness within the FL framework, owing to its training across disparate datasets with varying demographics and prevalence rates. This adaptability enhances the model’s generalization capacity, fostering resilience to variations in data.
@@
Recent research has suggested alternative methodologies for conducting federated averaging on non-independent and identically distributed (non-IID) data (Hong2022-yd, ; Tenison2022-uh, ; Xiao2020-sa, ). We used FedAvg as a baseline model for this study to establish preliminary feasibility. We chose our prediction labels based on ICD-10 codes associated with clinically important cardiovascular and metabolic conditions, prior demonstration of predictability using ECGs (Sun2022-nx, ), and sufficient representation across various hospital sites. However, we acknowledge the representative nature of our label selection and highlight the need for future experiments to explore if similar trends extend to other diagnostic categories. The use of ECG machines from a single manufacturer to capture all ECG data in this study presents a potential limitation, which may restrict the diversity of instrumentation and limit the generalizability of the study’s findings. However, the examination of ECG cohorts reveals notable variations in age, sex and disease prevalence rates across the 7 hospitals despite their geographical proximity and single healthcare provider (Table 2). These observed variations in clinical and demographic distributions among hospitals underscore the importance for adopting FL approaches for developing robust and equitable prediction models.
@@
In the DP experiment, the DP-SGD algorithm revealed a discernible correlation between the value of $\epsilon$ and the model performance. By constraining gradients and introducing controlled noise, DP-SGD mitigates the adversarial risk on the training process. Moreover, DP-SGD can safeguard against privacy breaches, including membership inference attacks, through strategic noise injection, hindering malicious actors from inferring sensitive information. As $\epsilon$ increases, indicating a decrease in privacy, a concurrent increase in model performance is observed—an anticipated outcome. While the augmented noise levels enhance privacy safeguards, they also diminish the model’s ability to discern subtle patterns in the ECG data. Consequently, the model’s performance deteriorates when $\epsilon$ values are reduced. This observation underscores the fundamental trade-off to be considered in applications of DP: decreasing privacy constraints may augment prediction performance, whereas increasing privacy can lead to compromised model performance. We observed that the trade-off between privacy and performance becomes more pronounced with diminishing $\epsilon$ values. Therefore, striking the right balance between preserving privacy and maintaining a high level of predictability remains a complex and ongoing challenge. One potential approach for determining the optimal value of $\epsilon$ in the context of DP might be employing Membership Inference Attacks (7958568, ) to evaluate the effectiveness of the privacy measures. Ponomareva et al. (Ponomareva2023-me, ) posits that maintaining an epsilon value below 10 provides a substantiated privacy assurance. Additionally, the process of computing noise to gradients can significantly increase the training time and resource requirements, which can be a practical concern in applications where real-time processing is critical. Therefore, practitioners must carefully consider these trade-offs and make informed decisions when implementing DP-SGD in their machine learning workflows. Nonetheless, the benefits of DP-SGD in terms of privacy preservation make it a compelling approach, with its limitations being areas for continued refinement in privacy-preserving machine learning.